OpenStudy (anonymous):

convert this equation from standerd form to slope-intercept form 3x-5y=25

OpenStudy (anonymous):

Solve for what? Two unknowns, only one equation. Solution is infinite set ... if that is what you really want. Or do you want this general linear equation written in slope-intercept form. Lucky for you that gives the answer to BOTH possible questions. 3x + 5y = 25 Subtract 3x ... 5y = –3x + 25 Divide by 5 . . . y = (–3/5)x + 5 or more neatly: y = –3x/5 + 5 slope-intercept form So the infinite set of solutions (that is if you really wanted it "solved") would be the set of ordered pairs {(a, –3a/5 +5) | a=arbitrary number} For example if a=10, we get (10,–3(10)/5+5) = (10, –1) as one of the infinity of solutions

OpenStudy (anonymous):

For multiple variables, you need as many unique equations as variables to solve for them. Because you have two variables and one equation, you can't solve for them numerically. This next thing I have to say is important: THERE ARE INFINITE NUMERICAL SOLUTIONS TO THIS EQUATION!!!! You may have notice that people have given you different answers, all of which that work. That is why. However, you can solve for them in terms of each other: 3x+5y=25 3x=25-5y 3x=5(5-y) x=(5/3)(5-y) Alternatively: 3x+5y=25 5y=25-3x y=5-(3/5)x So you can't actually get the numeric values because there isn't a single correct set of numeric values for X and Y, but you can solve them in terms of each other.

Nnesha (nnesha):

the equation in the question is 3x - 5y = 25 "Negative 5y" so you have to subtract 3x first  so left side positive 3x and negative 3x should cancel out and on right side we cannot combine 25 and -3x so -3x stay there